OpenStudy (anonymous):

What is the value of the expression? 1/4^3

OpenStudy (anonymous):

1/64

OpenStudy (whpalmer4):

@jpenrod okay, now explain how you got that

OpenStudy (anonymous):

Yes please explain 'cause I just don't understand.

OpenStudy (anonymous):

4x4=16 16x4= 64 1/64 simple

OpenStudy (anonymous):

So you just do 4x4x4=64 as the bottom part then just bring the 1 over to the top?

OpenStudy (whpalmer4):

you understand that 4^3 = 4*4*4, right?

OpenStudy (anonymous):

yes, i think i understand now. i should of just worked it out.

OpenStudy (anonymous):

What is the value of the expression? 2/4^-3 How do I do this one with the negative power?

OpenStudy (whpalmer4):

so 1/(4^3) = 1/(4*4*4) = 1/64 also, by the properties of exponents, (1/4)^3 also equals 1/64: (1/4)*(1/4)*(1/4) and finally, you might see this written as 4^-3, which is equivalent, because a^-n = 1/a^n

OpenStudy (whpalmer4):

good to be familiar with all the different ways you might see the same thing, because, well, you'll be seeing all of them sooner or later :-)

OpenStudy (whpalmer4):

do you have a few more of those? you can do them and I'll check your work

OpenStudy (anonymous):

A. 1/32 B. 1/6 C. 128 D. 256 OK So the answer is A.?

OpenStudy (whpalmer4):

what's the problem? :-)

OpenStudy (anonymous):

The one we just did, the second one.

OpenStudy (whpalmer4):

oh, I didn't see that you snuck that in there, sorry! :-)

OpenStudy (anonymous):

no its 128

OpenStudy (whpalmer4):

2/4^-3 well, let's first convert it to a positive power: if it's a negative power, we make it positive, and put it on the other side of the division line, so it becomes 2*4^3

OpenStudy (anonymous):

2/64 = 2*4^3 Is that right?

OpenStudy (whpalmer4):

2/64 = 2/4^3 or 2 * 4^-3 but are we talking about a different problem now? the second one you showed was 2/4^-3 or 2 ------ 4^-3 and because 4^-3 = 1/4^3 we can write it as 2 ------ (1/4^3) dividing a fraction by a fraction, you invert the fraction in the denominator and multiply: 2 * (4^3/1) = 2*4^3 = 2*64 = 128, like jpenrod said

OpenStudy (anonymous):

Sorry, OK The problem is What is the value of the expression? 2/4^-3 A. 1/32 B. 1/6 C. 128 D. 256

OpenStudy (whpalmer4):

right, so the answer is C. 128 do you understand how we got that answer?

OpenStudy (anonymous):

Yeah...I think so. :D

OpenStudy (whpalmer4):

well, let's try another and find out :-) what is the value of 4/3^-2?

OpenStudy (anonymous):

What about this one.. What value of k solves the equation? K-4 = 1/256 A. –64 B. 8 C. 4 D. 64

OpenStudy (whpalmer4):

let's do mine first :-)

OpenStudy (anonymous):

4/-18? I don't really think that's it. I'm bad with negatives

OpenStudy (whpalmer4):

remember: if you have a negative exponent, make it a positive exponent and put the quantity on the other side of the division bar. 4 4*3^2 ------ = ------- = ? 3^-2 1

OpenStudy (anonymous):

OK so 4/9? I know it somehow gets to 36, right?

OpenStudy (whpalmer4):

no, we had 4*3^2 ----- 1 3^2 = 9, right 4*9 ---- 1 = 36

OpenStudy (whpalmer4):

okay, your problem, is that actually supposed to be k^-4 ?

OpenStudy (whpalmer4):

because the answer to k-4 = 1/256 doesn't appear among the answer choices

OpenStudy (anonymous):

woops, sorry. k^-4

OpenStudy (whpalmer4):

right. so what is another way of writing k^-4?

OpenStudy (anonymous):

sorry, computer issues....IDK, is there

OpenStudy (anonymous):

@Yuba @whpalmer4

OpenStudy (whpalmer4):

k^-4 = 1/k^4, right? negative exponent becomes a positive exponent on the other side of the division bar

OpenStudy (whpalmer4):

so that gives us 1 1 --- = ----- k^4 256 how could you solve that?

OpenStudy (anonymous):

Find the common denominator?

OpenStudy (phi):

you could "flip" both sides

OpenStudy (phi):

or cross multiply

OpenStudy (anonymous):

what is K^4

OpenStudy (whpalmer4):

if those fractions are equal, and the numerators are identical, don't the denominators have to be equal?

OpenStudy (whpalmer4):

If I have 1 1 -- = -- x 4 doesn't x have to equal 4?

OpenStudy (whpalmer4):

that's the same as "flipping" both sides. cross-multiplying would give you the same thing: 4*1 = x*1 4 = x so here we have k^4 = 256 what is the value for k such that k*k*k*k = 256?

OpenStudy (anonymous):

4^4 = 256......right?

OpenStudy (anonymous):

So is the answer 4

OpenStudy (whpalmer4):

if k = 4, then (4)^-4 = 1/256 1/(4^4) = 1/256 1/(256) = 1/256 which is true. yes, the answer is 4.

OpenStudy (anonymous):

YAY!

OpenStudy (whpalmer4):

good job!